OpenStudy (anonymous):

Astronomers measure large distances in light-years. One light-year is the distance that light can travel in one year, or approximately 5.88 × 10^12 miles. Suppose a star is 9.8 × 10^1 light-years from Earth. In scientific notation, approximately how many miles is it?

OpenStudy (barrycarter):

9.8*10^1 light years * (5.88*10^12) miles/light year = 57.624*10^13 miles (by adding exponents for the 10^ stuff). Since scientific numbers must start with something between one and nine: 57.624*10^13 = 5.7624*10*10^13 = 5.7624*10^14